OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y = sqroot(x) Point: (64, 8)

OpenStudy (amistre64):

\[y-8=f'(64)(x-64)\]

OpenStudy (anonymous):

i need to use equation f(x) - f(a) /x-a to find slope first but im confused on what f(a) would be

OpenStudy (anonymous):

because f(x) would just be sqroot(x) right?

OpenStudy (amistre64):

right

OpenStudy (amistre64):

do you know deriviative rules yet?

OpenStudy (anonymous):

also i meant the lim as x approaches a for that equation...and slightly, derivatives are what we are just starting

OpenStudy (anonymous):

1/2SQROOT(x) 1/2(8) 1/16

OpenStudy (amistre64):

\[\frac{\sqrt{x}-\sqrt{a}}{x-a}\] \[\frac{\sqrt{x}-\sqrt{a}}{x-a}*\frac{\sqrt{x}+\sqrt{a}}{\sqrt{x}+\sqrt{a}}\] \[\frac{(x-a)}{(x-a)(\sqrt{x}+\sqrt{a})}\] \[\frac{1}{\sqrt{x}+\sqrt{a}}\]

OpenStudy (amistre64):

when x = a we have the slope at x

OpenStudy (anonymous):

im a little confused ha...so a would equal x?

OpenStudy (amistre64):

\[\lim_{a->x}\frac{1}{\sqrt{x}+\sqrt{a}}=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

y = (x)^1/2 ,y '=1/2sqrtx, using the point (64, 8) at x=64, y '=1/2sqrt64=1/16 y-64=(1/16)(x-8) y= (x/16)+4

OpenStudy (amistre64):

yes, when a = x is what we get when as a gets to x

OpenStudy (anonymous):

so it would be sqroot(a)

OpenStudy (amistre64):

not if you follow the algebra :)

OpenStudy (amistre64):

1/2sqrt(a); but it makes more sense in the problem to say that a approaches 64 until a = 64

OpenStudy (amistre64):

and since x=64 ....

OpenStudy (anonymous):

im lost haha...i dont understand

OpenStudy (anonymous):

would a = 64? 2(sqroot(8)) ?

OpenStudy (anonymous):

er 2(sqroot(64))

OpenStudy (anonymous):

thanks i worked out the problem